# Emergence of a second law of thermodynamics in isolated quantum systems

### v1

@@ II.3 Replacing projective observables with POVMs
$S_{\mathrm{Sh}}^{M}[\rho]=-\sum_{i=1}^{r}p_{i}\log p_{i},\text{ with }p_{i}:=\textnormal{Tr}[E_{i}\rho],$ (65)
defined analogously to the Shannon entropy of the probability vector $\vec{p}=(p_{1},\dots,p_{r})$ relative to the POVM $M$. The physical interpretation remains the same for this entropy as for the notion defined relative to an observable; the entropy $S_{\mathrm{Sh}}^{M}[\rho]$ is a measure of the information relative to $M$ that one gains by performing a measurement. Similarly, the observational entropy can also be generalized using
$S_{\mathrm{Obs}}^{M}[\rho]=-\sum_{i=1}^{r}p_{i}\log\frac{p_{i}}{V_{i}},\text{ with }p_{i}:=\textnormal{Tr}[E_{i}\rho],\,V_{i}=\textnormal{Tr}[E_{i}],$ (66)
@@
In the main text we looked at how the entropy of $\rho(t)$ relative to an observable $O$ compares to the entropy of $\omega$ (relative to the same observable $O$). We wish to generalize these statements to the case where the entropy is now defined relative to the POVM $M$ in question. Since our main results Theorem 1, Prop. 2 and Cor. III all follow from the asymptotic closeness of the population vector $\vec{p}(t)$ to the equilibrium populations $\vec{p}_{\infty}$, what must be shown is that for a more general POVM $M$, the two population vectors are also close for most times. It turns out that Lemma II.1 also generalizes to the case of POVMs:
@@
The proof is similar to that of Lemma II.1, with the difference that while for projectors we can use the fact that $\Pi_{i}^{2}=\Pi_{i}$, this is generally not the case for a POVM: $E_{i}^{2}\neq E_{i}$. Therefore, we have to replace Eq. (42) with
@@
again making use of the completeness relation $\sum_{i=1}^{r}E_{i}=\mathds{1}.$
As a consequence, the resulting inequalities from Theorem 1 for the Shannon entropy $S_{\mathrm{Sh}}^{M}$ and from Prop. 2 for the observational entropy $S_{\mathrm{Obs}}^{M}$ also follow for the POVM $M$. For the Shannon entropy bound, one can directly use Lemma II.3 instead of Lemma II.1 to prove the statement. On the other hand, for the observational entropy, the proof has to be modified such that the coarse-grained state is defined with respect to the general POVM elements in $M$, instead of using the definition with respect to the orthonormal projectors that we have for an observable $O$. To be precise, for some quantum state $\rho$ we define the coarse-grained version as
$\rho^{\mathrm{cg}}=\sum_{i=1}^{r}\textnormal{Tr}[E_{i}\rho]\frac{E_{i}}{V_{i}},$ (71)
where $V_{i}=\textnormal{Tr}[E_{i}]$ as before. This also gives us a well-defined quantum state [38], and the remaining proof of Prop. 2 is exactly analogous as for the case of observables. A useful consequence of this generalization is that with a POVM one can model measurements which are coarse-grained in the sense that they can be used to model measurements which do not perfectly distinguish some states.
Appendix C Details of the numerical model used
Here, we describe in more detail the model used to generate the numerical results in Figs. 1, 2 and 3. The model used is given in Ref. [41] with the Hamiltonian
$H=\sum_{i=1}^{N}g\sigma_{x}^{(i)}+\sum_{i=2}^{N-1}h\sigma_{z}^{(i)}+(h-J)\left(\sigma_{z}^{(1)}+\sigma_{z}^{(N)}\right)+\sum_{i=1}^{N-1}J\sigma_{z}^{(i)}\sigma_{z}^{(i+1)},$ (72)

### v2

@@ Appendix C Replacing projective observables with POVMs
$S_{\mathrm{Sh}}^{M}[\rho]=-\sum_{i=1}^{r}p_{i}\log p_{i},\quad p_{i}:=\textnormal{Tr}[E_{i}\rho],$ (72)
defined analogously to the Shannon entropy of the probability vector $\vec{p}=(p_{1},\dots,p_{r})$ relative to the POVM $M$. The physical interpretation remains the same for this entropy as for the notion defined relative to an observable; the entropy $S_{\mathrm{Sh}}^{M}[\rho]$ is a measure of the information relative to $M$ that one gains by performing a measurement. It is important to note, however, that some of the entropy may when using POVMs may not be due to an intrinsic uncertainty coming from the underlying quantum state as is the case for PVMs, but some of the uncertainty comes from the coarse graining of the POVM. An examplary case is if $M=(E_{1},E_{2})$ and $E_{1}=E_{2}=\mathds{1}/2,$ the Shannon observable entropy is $\log 2$, not due to the underlying state but due to the randomness coming from the POVM. Similarly, the observational entropy can also be generalized using
$S_{\mathrm{Obs}}^{M}[\rho]=-\sum_{i=1}^{r}p_{i}\log\frac{p_{i}}{V_{i}},\quad p_{i}:=\textnormal{Tr}[E_{i}\rho],\,V_{i}=\textnormal{Tr}[E_{i}],$ (73)
@@
In the main text we looked at how the entropy of $\rho(t)$ relative to an observable $O$ compares to the entropy of $\omega$ (relative to the same observable $O$). We wish to generalize these statements to the case where the entropy is now defined relative to the POVM $M$ in question. Since our main results Theorem 2, Proposition 3 and Corollary III all follow from the asymptotic closeness of the population vector $\vec{p}(t)$ to the equilibrium populations $\vec{p}_{\infty}$, what must be shown is that for a more general POVM $M$, the two population vectors are also close for most times. It turns out that Lemma I.1 also generalizes to the case of POVMs:
@@
The proof is similar to that of Lemma I.1, with the difference that while for projectors we can use the fact that $\Pi_{i}^{2}=\Pi_{i}$, this is generally not the case for a POVM: $E_{i}^{2}\neq E_{i}$. Therefore, we have to replace Eq. (23) with
@@
again using the completeness relation $\sum_{i=1}^{r}E_{i}=\mathds{1}.$
As a consequence, the resulting inequalities from Theorem 2 for the Shannon entropy $S_{\mathrm{Sh}}^{M}$ and from Proposition 3 for the observational entropy $S_{\mathrm{Obs}}^{M}$ also follow for the POVM $M$. For the Shannon entropy bound, one can directly use Lemma C instead of Lemma I.1 to prove the statement. On the other hand, for the observational entropy, the proof of Lemma 3 has to be modified to also hold for a general POVM $M$, instead of using the definition with respect to the orthonormal projectors that we have for an observable $O$. The proof goes through completely analogously simply by replacing eq. (60) with
$\vec{p}_{\rho}=\sum_{i=1}^{r}\textnormal{Tr}[E_{i}\rho]|{i}\rangle\!\langle{i}|,$ (78)
thus generalizing both Lemma 3 and Proposition 3 to the case of POVMs. A useful consequence of this generalization is that with a POVM one can model measurements which are coarse-grained in the sense that they can be used to model measurements which do not perfectly distinguish some states.
Appendix D Details of the numerical model used
Here, we describe in more detail the model used to generate the numerical results in Figs. 1, 3 and 2. The model used is given in Ref. [43] with the Hamiltonian
$H=\sum_{i=1}^{N}g\sigma_{x}^{(i)}$ $+\sum_{i=2}^{N-1}h\sigma_{z}^{(i)}+(h-J)\left(\sigma_{z}^{(1)}+\sigma_{z}^{(N)}\right)$
$+\sum_{i=1}^{N-1}J\sigma_{z}^{(i)}\sigma_{z}^{(i+1)},$ (79)